# Hermitian and non-Hermitian topology in active matter

### v1

@@ 4.1.1 Topology in passive fluid
While early studies of topological physics were motivated by electronic band structures, recent studies have also revealed the existence of topological edge modes in classical systems. One of the platforms for exploring such classical topological waves is fluid, such as air and water in engineered structures [130, 132, 133, 247]. In topological fluids, one can observe the robust propagation of density waves, i.e., sound, along the boundary of the system, which is analogous to the edge current in topological insulators. Below we review the theoretical basis of topological fluids.
@@
To further simplify the equation and make its connection to the Schrödinger equation clear, it is useful to assume that the viscosity and heat flow are negligible. The equations (4.1.1) and (77) can then be converted into
@@
where $\phi$ is the velocity potential that satisfies $\mathbf{v}=-\nabla\phi$ [130]. After the Fourier transformation, the time derivative $\partial_{t}$ is replaced by $i\omega$, where $\omega$ represents the frequency of a standing wave. By assuming $|\mathbf{v}/c|\ll 1$ and thus ignoring the higher-order terms, one finally obtains the following Schrödinger-like equation [130]:
$\frac{\omega^{2}}{c^{2}}\Psi=[(-i\nabla+\mathbf{A}_{\rm eff})^{2}+V]\Psi$ (79)
@@
$\mathbf{A}_{\rm eff}$ $=$ $-\omega\mathbf{v}_{0}/c^{2},$ (80)
$V$ $=$ $-|\nabla\ln\rho|^{2}/4-\Delta\ln\rho/2+\omega^{2}/c^{2}.$ (81)
@@
The derived Schrödinger-like equation tells us that the steady-state flow plays a role of the effective vector potential, whose vorticity corresponds to the effective (local) magnetic flux that is necessary to realize a counterpart of the quantum (anomalous) Hall effect. Therefore, if the steady-state flow in a phononic system imitates the effective vector potential in quantum (anomalous) Hall systems, such a system should exhibit topological boundary waves. One straightforward way of constructing a topological fluid is globally rotating fluid by using circulators or external forces to a container of fluid [250, 251]. The global circular motion of fluid induces a net effective magnetic field, and thus the system constructs a fluidic counterpart of the quantum Hall effect. This correspondence can also be understood from the rotating coordinate system; if we consider the rotating coordinate that cancels the rotational motion of fluid, the fluid feels a Coriolis force. The Coriolis force acts perpendicularly to the direction of flow, which is the same as a (homogeneous) magnetic force. We can thus again assume that the fluid feels the effective magnetic field, and such effective magnetic force leads to the phononic counterpart of the quantum Hall effect.
@@
Topological fluid has been also realized without a net effective magnetic field by using an analogy to the quantum anomalous Hall effect. We note that the quantum anomalous Hall effect also requires the breakdown of the time-reversal symmetry. Therefore, purely passive fluids cannot construct a fluid counterpart of the quantum anomalous Hall effect, and one must utilize external driving devices such as circulators [248]. One of the early studies [130] utilized periodically aligned circulators, where circulators are placed on each lattice point of a triangular lattice,
@@
Topological fluids are not restricted to counterparts of the quantum (anomalous) Hall effect. Without breaking the time-reversal symmetry (i.e., no external drivings), one can realize acoustic quantum spin Hall systems. Reference [247] has realized such a quantum spin Hall system, by using a honeycomb-lattice structure and assuming left and right-rotational modes of sound waves as effective spins. One can also realize higher-order topological insulators of fluids. Since the width of a channel corresponds to the strength of a hopping, one can realize staggered hoppings that are akin to the SSH and BBH models. By using such a technique, topological corner modes have been proposed and observed in acoustic metamaterials [249, 252].
@@
To derive the sound master equation (78), we have ignored dissipative terms originating from the viscosity; however, such dissipative terms can in general make the effective Hamiltonian non-Hermitian. Some papers also explore the combination of such non-Hermiticity and topology of fluids [253, 254]. In particular, the non-Hermitian skin effect has been experimentally realized [192] by using feedback controls.
@@
Some of the theoretical proposals of topological phononics are realized in experiments of acoustic metamaterials. In particular, chiral edge modes using circulation devices are observed in a metamaterial with a honeycomb-lattice structure [248]. Another intriguing example of applications of topological phononics can be seen in geophysics [250]. Around the equator, unidirectional waves called the Yanai wave and Kelvin wave have been observed. A recent study has revealed that those equatorial waves can be a consequence of the nontrivial topology of the fluid (water or atmosphere) on a sphere. As discussed above, the Coriolis force can be regarded as an effective magnetic force, and its direction changes when fluid crosses the equator. Therefore, the seawater on the north and the south hemispheres exhibits opposite Chern numbers, and the topological modes appear at the boundary. i.e., the equator.
@@
In the last part of this section, we mention the possible breakdown of the bulk-boundary correspondence in continuum systems including fluids. Unlike condensed matter systems with discrete translation symmetries, fluids often exhibit continuum translation symmetries. In such a case, one cannot consider a finite-size wavenumber space (i.e. a Brillouin zone) but should assume an infinite wavenumber space, where the wavenumber can go to infinity. Mathematically, the finite and infinite wavenumber spaces can be distinguished by their compactness. Thus, the conventional topological band theory can be broken in continuum systems, as theoretically predicted in certain phononic systems by numerically confirming the inconsistency between the topological invariant and the number of gapless modes [255]. The key to recover the bulk-boundary correspondence is the compactification of the wavenumber space; one should construct an effective Hamiltonian $H(\mathbf{k})$ that exhibits the asymptotically same behavior in the limit of $|\mathbf{k}|\rightarrow\infty$ independently of the direction of $\mathbf{k}$. This compactification can be done by utilizing, e.g., odd-viscosity [251].
@@ 4.1.2 Other topological phononics
Non-Hermiticity is also ubiquitous in classical systems because they are typically open systems exchanging energies with environments, and thus their fundamental equations include dissipative terms. In mechanical lattices, friction is a major origin of the non-Hermiticity, while its interplay with topological edge modes is largely unexplored. Another possibility to introduce the non-Hermitian interaction is a nonreciprocal force generated by elaborate devices such as fans attached to a mass point whose direction depends on the distance between two mass points [261]. By introducing nonreciprocal forces into mechanical lattices, some previous studies [9, 262] have investigated the non-Hermitian skin effect in classical systems.
@@ 4.2.1 Classical analogs of quantum Hall effect
As in passive fluid, one can construct an analogy between the linearized hydrodynamics of self-propelled particles and the Schrödinger equation. Recent studies [30, 190, 251, 263] have proposed active-matter counterparts of topological insulators based on such linearized hydrodynamics. Starting from the Toner-Tu equation, one can obtain the following linearized equation,
@@
where $\delta\rho=\rho-\rho_{0}$ and $\delta\mathbf{v}=\mathbf{v}-\mathbf{v}_{0}$ are the fluctuations from the steady state density $\rho_{0}$ and velocity field $\mathbf{v}_{0}$, respectively. Here, we ignore the advection terms including $\lambda_{2},\lambda_{3}$ since their effect can be included by renormalization of $\lambda$. We also ignore the diffusion terms to focus on the Hermitian topology. If we consider the short-term dynamics of fluctuations compared to the inverse of the decaying rate, we can assume that such dynamics can be captured by a Hermitian matrix without dissipative terms, though nonreciprocity or exceptional points can break such an assumption as we will see later. Furthermore, by assuming that the speed of active particles is $v_{0}=\sqrt{\alpha/\beta}$, the first term in the right-hand side of the second equation exhibits no $\alpha$ dependence. In this linearized hydrodynamics, the effect of the activity appears in the additional parameter $\lambda$ and the higher-order term $-2\beta(\mathbf{v}_{\rm ss}\cdot\delta\mathbf{v})\mathbf{v}_{\rm ss}$ of the steady flow. However, if we further assume that the steady-state vector field $\mathbf{v}_{\rm ss}$ is small enough compared to the effective sound speed $c$ (cf. Eq. (5)), one can derive the Schrödinger-like equation
@@
as in passive fluids. In this equation, $\mathbf{V}_{0}=\omega(\lambda+1)\mathbf{v}_{0}/2$ is the effective vector potential that is proportional to the steady-state flow $\mathbf{v}_{0}$. The advantage of the active system is that the steady-state flow appears spontaneously, i.e., without using external drivings.
@@
To construct active-matter counterparts of the quantum Hall effect, one needs to realize the effective magnetic field. Unlike passive fluids, effective vector potentials can be created by the spontaneous flow of active particles. However, if only a simple structure is available, the effective magnetic fields must be globally canceled, and analogs of the quantum Hall effect are prohibited [264]. To avoid the cancellation of local effective magnetic fields, a pioneering work of topological active matter [30] has proposed to utilize channels aligned on a Lieb-lattice structure that has defective sites [Fig. 18(a)]. In the Lieb-lattice channels, clockwise and anticlockwise flows are alternately generated by the spontaneous motion of active matter, and they create local magnetic fields with opposite signs. If we consider a simple square lattice, the numbers of clockwise and anticlockwise flows are the same, which implies the cancellation of the effective magnetic field. Meanwhile, the defects in the Lieb lattice lead to the imbalance of the number of the clockwise and anticlockwise vortices, and thus the net effective magnetic field remains nonzero. As a consequence of the nontrivial topology due to the nonzero effective magnetic field, localized boundary modes are observed in the numerical calculation of the eigenvectors of the effective Hamiltonian of the active metamaterial. To experimentally observe such localized boundary modes, one should oscillate the active fluid at the source point, and confirm the existence of the chiral edge wave. Such a chiral edge mode has been also confirmed in the numerical simulation of the Lieb-lattice model.
@@
The counterpart of the quantum Hall effect can also be constructed from active matter on curved surfaces [263, 265]. As in a Lieb-lattice model, spontaneous flows of active matter induce effective magnetic fields without the help of, e.g. the external rotation (cf. the self-rotation of a sphere). For example, active matter on a sphere goes around and creates north and south poles and hemispheres reflecting on the direction of rotation. Then, the localized modes can be observed at the boundary of hemispheres, which is analogous to the Yanai and Kelvin waves on the Earth. One surprising point is that the spontaneous flow is stabilized by the activity and interactions of active particles, and thus the topological boundary modes can emerge independently of an initial condition. Topological boundary modes can also emerge on more complex surfaces, such as a torus and a helicoidal sphere [265]. As we can infer from the physics of the Coriolis force, the sign of the effective magnetic force and the Chern number are changed by the sign of the change rate of the radius.
@@
Chiral active matter is also useful to realize topological active matter [190, 251] because its chirality naturally leads to the effective Lorentz force. Since chiral active matter often shows self-rotation, when it goes to, e.g., the left direction at a certain moment, it tends to move in the upper or lower direction at the next moment. This implies that chiral active matter feels a rotational force that is perpendicular to the velocity of particles, which is analogous to the Lorentz force. The effective Lorentz force can appear as the term $\omega_{0}\nabla\times\mathbf{v}$ in the hydrodynamic equations. By linearizing the hydrodynamics of chiral active matter, we obtain the following equation,
@@
The subtle point in the bulk-boundary correspondence in fluids mentioned above can be also seen in active hydrodynamics. Specifically, without the odd-viscosity term, the gapless boundary modes disappear [265] even if the bulk bands have nontrivial topology. Fortunately, the odd viscosity can ubiquitously emerge in chiral active matter as discussed in Sec. 2, and thus we can expect the conventional bulk-boundary correspondence between the Chern number and edge modes in active fluid.
Topological active matter utilizing chirality seems to be more amenable to experimental realizations since there is no need to prepare engineered structures. Indeed, some studies have utilized artificial chiral active matter, such as robotic rotators [266] and vibrated granular gears [267], as well as active spinners [269, 270]. The robotic rotators have an asymmetric structure and self-vibration using motors, leading to their rotational movements. After preparing the collection of such robotic rotators, one can observe the localization of the distribution of the rotators as discussed in Ref. [190]. Vibrated granular gears have also asymmetric legs, and external vibration leads to their active motion. If one puts a passive cargo in the collection of the granular gears, the robust transport of the cargo can be observed along the edge of the system, which is analogous to the topological boundary modes discussed in Ref. [251].
Topological chiral active matter is also realizable in biological systems. An experimental study [268] shows the chiral collective motion of neural progenitor cells. Specifically, when confining such cells in a wide channel, a chiral boundary flow of cells can be observed. Boundary modes are also found in power spectra of fluctuations of the flow field, which directly confirms the existence of topological edge modes in the linearized hydrodynamics. The neural progenitor cells are self-propelled rods that involve nematic tensors when described in the hydrodynamic equations. One can derive linearized equations similar to those obtained from active hydrodynamics of polar active matter and numerically confirm its nonzero Chern number. The importance of odd-viscosity terms in the bulk-boundary correspondence in neural progenitor cells has also been discussed [268].

### v2

@@ 5.1.1 Topology in passive fluid
While early studies of topological physics were motivated by electronic band structures, recent studies have also revealed the existence of topological edge modes in classical systems. One of the platforms for exploring such classical topological waves is fluid, such as air and water in engineered structures [133, 135, 136, 255]. In topological fluids, one can observe the robust propagation of density waves, i.e., sound, along the boundary of the system, which is analogous to the edge current in topological insulators (cf. Fig. 4(a)). Below we review the theoretical basis of topological fluids, whose extension is also utilized in theoretical proposals of topological active matter.
@@
To further simplify the equation and make its connection to the Schrödinger equation clear, it is useful to assume that the viscosity and heat flow are negligible. Applying $(\partial_{t}+\mathbf{v}_{0}\cdot\nabla)$ to the left-hand side of Eq. (5.1.1) and rewriting $(\partial_{t}+\mathbf{v}_{0}\cdot\nabla)\delta\mathbf{v}$ by using Eq. (78) without the viscosity term $D_{T}\nabla^{2}\delta\mathbf{v}$ and the term related to heat flow $(\delta\mathbf{v}\cdot\nabla)\mathbf{v}_{0}$, one obtains
$(\partial_{t}+\mathbf{v}_{0}\cdot\nabla)^{2}\delta\rho=\rho_{0}p_{0}\Delta\delta\rho.$ (79)
Then, one rewrites $p_{0}$ in terms of the sound velocity $c$ as $p_{0}=c^{2}/\rho_{0}$ and assumes $|\mathbf{v}/c|\ll 1$. By considering the Fourier component $\delta\rho(t)=e^{-i\omega t}\delta\rho^{\prime}$, one can obtain the following Schrödinger-like equation:
$\frac{\omega^{2}}{c^{2}}\delta\rho=(-i\nabla+\mathbf{A}_{\rm eff})^{2}\delta\rho.$ (80)
This equation is reminiscent of the conventional Schrödinger equation (cf. Eqs. (22) and (27)) without a scaler potential. Comparing this equation with the Schrödinger equation, one can assume $\mathbf{A}_{\rm eff}=\omega\mathbf{v}_{0}/c^{2}$ as an effective vector potential.
One can also derive another Schrödinger-like equation via the following sound master equation [133]
@@
where $\phi$ is the velocity potential that satisfies $\mathbf{v}=-\nabla\phi$ and $c$ corresponds to the sound velocity. This sound master equation is derived from the Navier-Stokes equation under the same assumption (the viscosity and heat flow are negligible) as in Eq. (79). After the Fourier transformation, the time derivative $\partial_{t}$ is replaced by $i\omega$, where $\omega$ represents the frequency of a standing wave. By assuming $|\mathbf{v}/c|\ll 1$ and thus ignoring the higher-order terms as in Eq. (79), one finally obtains the following Schrödinger-like equation [133]:
$\frac{\omega^{2}}{c^{2}}\Psi=[(-i\nabla+\mathbf{A}^{\prime}_{\rm eff})^{2}+V]\Psi$ (82)
@@
$\mathbf{A}^{\prime}_{\rm eff}$ $=$ $-\omega\mathbf{v}_{0}/c^{2},$ (83)
$V$ $=$ $-|\nabla\ln\rho|^{2}/4-\Delta\ln\rho/2.$ (84)
@@
The derived Schrödinger-like equations tell us that the steady-state flow plays a role of the effective vector potential, whose vorticity corresponds to the effective (local) magnetic flux that is necessary to realize a counterpart of the quantum (anomalous) Hall effect. Therefore, if the steady-state flow in a phononic system imitates the effective vector potential in quantum (anomalous) Hall systems, such a system should exhibit topological boundary waves. One straightforward way of constructing a topological fluid is globally rotating fluid by using circulators or external forces to a container of fluid [258, 259]. The global circular motion of fluid induces a net effective magnetic field. Since the nonzero net magnetic field is a key ingredient of the quantum Hall effect (cf. Sec. 3.2.1), the rotating system constructs a fluidic counterpart of the quantum Hall effect. This correspondence can also be understood from the rotating coordinate system; if we consider the rotating coordinate that cancels the rotational motion of fluid, the fluid feels a Coriolis force. The Coriolis force acts perpendicularly to the direction of flow, which is the same as a (homogeneous) magnetic force. We can thus again assume that the fluid feels the effective magnetic field, and such effective magnetic force leads to the phononic counterpart of the quantum Hall effect.
@@
Topological fluid has been also realized without a net effective magnetic field by using an analogy to the quantum anomalous Hall effect. We note that the quantum anomalous Hall effect also requires the breakdown of the time-reversal symmetry. Therefore, purely passive fluids cannot construct a fluid counterpart of the quantum anomalous Hall effect, and one must utilize external driving devices such as circulators [256]. One of the early studies [133] utilized periodically aligned circulators, where circulators are placed on each lattice point of a triangular lattice. Then, the engineered vector potential imitates that of Haldane’s honeycomb model [144] in Fig. 5(a) and thus constructs the counterpart of the quantum anomalous Hall effect.
@@
Topological fluids are not restricted to counterparts of the quantum (anomalous) Hall effect. Without breaking the time-reversal symmetry (i.e., no external drivings), one can realize acoustic quantum spin Hall systems. Reference [255] has realized such a quantum spin Hall system, by using a honeycomb-lattice structure and assuming left and right-rotational modes of sound waves as effective spins, which looks like the Kane-Mele model discussed in Fig. 6. One can also realize higher-order topological insulators of fluids. Since the width of a channel corresponds to the strength of a hopping, one can realize staggered hoppings that are akin to the SSH and BBH models (cf. Fig. 8). By using such a technique, topological corner modes have been proposed and observed in acoustic metamaterials [257, 260] (cf. Fig. 17(b)).
@@
To derive the sound master equation (81), we have ignored dissipative terms originating from the viscosity; however, such dissipative terms can in general make the effective Hamiltonian non-Hermitian. Some papers also explore the combination of such non-Hermiticity and topology of fluids [261, 262]. In particular, the non-Hermitian skin effect (cf. Sec. 4.3) has been experimentally realized [200] by using feedback controls.
@@
Some of the theoretical proposals of topological phononics are realized in experiments of acoustic metamaterials. In particular, chiral edge modes using circulation devices are observed in a metamaterial with a honeycomb-lattice structure [256] in Fig. 17(a). Another intriguing example of applications of topological phononics can be seen in geophysics [258]. Around the equator, unidirectional waves called the Yanai wave and Kelvin wave have been observed. A recent study has revealed that those equatorial waves can be a consequence of the nontrivial topology of the fluid (water or atmosphere) on a sphere. As discussed above, the Coriolis force can be regarded as an effective magnetic force, and its direction changes when fluid crosses the equator. Therefore, the seawater on the north and the south hemispheres exhibits opposite Chern numbers, and the topological modes appear at the boundary. i.e., the equator.
@@
In the last part of this section, we mention the possible breakdown of the bulk-boundary correspondence in continuum systems including fluids. Unlike condensed matter systems with discrete translation symmetries (defined in Eq. (24)), fluids often exhibit continuum translation symmetries if there are no periodic structures that confine the fluids and thus the system is spatially homogeneous. In such a case, one cannot consider a finite-size wavenumber space (i.e. a Brillouin zone) but should assume an infinite wavenumber space, where the wavenumber can go to infinity. Mathematically, the finite and infinite wavenumber spaces can be distinguished by their compactness. Thus, the conventional topological band theory can be broken in continuum systems, as theoretically predicted in certain phononic systems by numerically confirming the inconsistency between the topological invariant and the number of gapless modes [263]. The key to recover the bulk-boundary correspondence is the compactification of the wavenumber space; one should construct an effective Hamiltonian $H(\mathbf{k})$ that exhibits the asymptotically same behavior in the limit of $|\mathbf{k}|\rightarrow\infty$ independently of the direction of $\mathbf{k}$. This compactification can be done by utilizing, e.g., odd-viscosity [259].
@@ 5.1.2 Other topological phononics
Non-Hermiticity is also ubiquitous in classical systems because they are typically open systems exchanging energies with environments, and thus their fundamental equations include dissipative terms. In mechanical lattices, friction is a major origin of the non-Hermiticity, while its interplay with topological edge modes is largely unexplored. Another possibility to introduce the non-Hermitian interaction is a nonreciprocal force generated by elaborate devices such as fans attached to a mass point whose direction depends on the distance between two mass points [269]. By introducing nonreciprocal forces into mechanical lattices, some previous studies [9, 192, 193] have investigated the non-Hermitian skin effect in classical systems.
@@ 5.2.1 Classical analogs of quantum Hall effect
As in passive fluid, one can construct an analogy between the linearized hydrodynamics of self-propelled particles and the Schrödinger equation. Recent studies [28, 198, 259, 270] have proposed active-matter counterparts of topological insulators based on such linearized hydrodynamics. Starting from the Toner-Tu equations (Eqs. (5) and (5)), one can obtain the following linearized equation,
@@
where $\delta\rho=\rho-\rho_{0}$ and $\delta\mathbf{v}=\mathbf{v}-\mathbf{v}_{0}$ are the fluctuations from the steady state density $\rho_{0}$ and velocity field $\mathbf{v}_{0}$, respectively. Here, we ignore the advection terms including $\lambda_{2},\lambda_{3}$ since their effect can be included by renormalization of $\lambda$ [115]. We also ignore the diffusion terms to focus on the Hermitian topology. If we consider the short-term dynamics of fluctuations compared to the inverse of the decaying rate, we can assume that such dynamics can be captured by a Hermitian matrix without dissipative terms, though nonreciprocity or exceptional points can break such an assumption as we will see later. Furthermore, by assuming that the speed of active particles is $v_{0}=\sqrt{\alpha/\beta}$, the first term in the right-hand side of the second equation exhibits no $\alpha$ dependence. In this linearized hydrodynamics, the effect of the activity appears in the additional parameter $\lambda$ and the higher-order term $2\beta(\mathbf{v}_{\rm ss}\cdot\delta\mathbf{v})\mathbf{v}_{\rm ss}$ of the steady flow. However, if we further assume that the steady-state vector field $\mathbf{v}_{\rm ss}$ is small enough compared to the effective sound speed $c$ (cf. Eq. (5)), one can derive the Schrödinger-like equation
@@
as in passive fluids (cf. Eq. (80)). In this equation, $\mathbf{V}_{0}=\omega(\lambda+1)\mathbf{v}_{0}/2$ is the effective vector potential that is proportional to the steady-state flow $\mathbf{v}_{0}$. The advantage of the active system is that the steady-state flow appears spontaneously, i.e., without using external drivings.
@@
To construct active-matter counterparts of the quantum Hall effect, one needs to realize the effective magnetic field as we have discussed in Sec. 3.2.1 for hard condensed matter and in Sec. 5.1.1 for passive fluids. Unlike passive fluids, effective vector potentials can be created by the spontaneous flow of active particles. However, if only a simple structure is available, the effective magnetic fields must be globally canceled, and analogs of the quantum Hall effect are prohibited [271]. To avoid the cancellation of local effective magnetic fields, a pioneering work of topological active matter [28] has proposed to utilize channels aligned on a Lieb-lattice structure that has defective sites [Fig. 18(a)]. In the Lieb-lattice channels, clockwise and anticlockwise flows are alternately generated by the spontaneous motion of active matter, and they create local magnetic fields with opposite signs. If we consider a simple square lattice, the numbers of clockwise and anticlockwise flows are the same, which implies the cancellation of the effective magnetic field. Meanwhile, the defects in the Lieb lattice lead to the imbalance of the number of the clockwise and anticlockwise vortices, and thus the net effective magnetic field remains nonzero. As a consequence of the nontrivial topology due to the nonzero effective magnetic field, localized boundary modes are observed in the numerical calculation of the eigenvectors of the effective Hamiltonian of the active metamaterial. To experimentally observe such localized boundary modes, one should oscillate the active fluid at the source point, and confirm the existence of the chiral edge wave. Such a chiral edge mode has been also confirmed in the numerical simulation of the Lieb-lattice model.
@@
The counterpart of the quantum Hall effect can also be constructed from active matter on curved surfaces [270, 272]. As in a Lieb-lattice model, spontaneous flows of active matter induce effective magnetic fields without the help of, e.g. the external rotation (cf. the self-rotation of a sphere). For example, active matter on a sphere goes around and creates north and south poles and hemispheres reflecting on the direction of rotation. Then, the localized modes can be observed at the boundary of hemispheres [Fig. 18(b)], which is analogous to the Yanai and Kelvin waves on the Earth discussed in Sec. 5.1.1. One surprising point is that the spontaneous flow is stabilized by the activity and interactions of active particles, and thus the topological boundary modes can emerge independently of an initial condition. Topological boundary modes can also emerge on more complex surfaces, such as a torus and a helicoidal sphere [272]. When such a complex surface is described as $(x,y,z)=(r(\theta)\cos\phi,r(\theta)\sin\phi,\zeta(\theta))$, the direction of the effective magnetic force depends on the sign of $dr/d\theta$ as the direction of the Coriolis force differs in the north and south hemisphere. Therefore, the Chern number is also changed by $dr/d\theta$, and topological modes appear at the boundary.
@@
Chiral active matter, i.e., active matter exhibiting rotating motions (cf. Sec. 2.3.3) is also useful to realize topological active matter [198, 259] because its chirality naturally leads to the effective Lorentz force [Fig. 19]. Since chiral active matter often shows self-rotation, when it goes to, e.g., the left direction at a certain moment, it tends to move in the upper or lower direction at the next moment. This implies that chiral active matter feels a rotational force that is perpendicular to the velocity of particles, which is analogous to the Lorentz force. The effective Lorentz force can appear as the term $\omega_{0}\nabla\times\mathbf{v}$ in the hydrodynamic equations. By linearizing the hydrodynamics of chiral active matter, we obtain the following equation,
@@
The subtle point in the bulk-boundary correspondence in fluids mentioned above (Sec. 5.1.1) can be also seen in active hydrodynamics. Specifically, without the odd-viscosity term, the gapless boundary modes disappear [272] even if the bulk bands have nontrivial topology. Fortunately, the odd viscosity can ubiquitously emerge in chiral active matter as discussed in Sec. 2, and thus we can expect the conventional bulk-boundary correspondence between the Chern number and edge modes in active fluid.
Topological active matter utilizing chirality seems to be more amenable to experimental realizations since there is no need to prepare engineered structures. Indeed, some studies have utilized artificial chiral active matter, such as robotic rotators [273] and vibrated granular gears [274], as well as active spinners [276, 277], which are introduced in Sec. 2.1.4. The robotic rotators have an asymmetric structure and self-vibration using motors, leading to their rotational movements. After preparing the collection of such robotic rotators, one can observe the localization of the distribution of the rotators as discussed in Ref. [198]. Vibrated granular gears have also asymmetric legs, and external vibration leads to their active motion. If one puts a passive cargo in the collection of the granular gears, the robust transport of the cargo can be observed along the edge of the system, which is analogous to the topological boundary modes discussed in Ref. [259].
Since chirality is abundant in biological active matter as is discussed in Secs. 2.1.2 and 2.1.3, topological chiral active matter is also realizable in biological systems. An experimental study [275] shows the chiral collective motion of neural progenitor cells. Specifically, when confining such cells in a wide channel, a chiral boundary flow of cells can be observed. Boundary modes are also found in power spectra of fluctuations of the flow field, which directly confirms the existence of topological edge modes in the linearized hydrodynamics. The neural progenitor cells are self-propelled rods that involve nematic tensors when described in the hydrodynamic equations; their hydrodynamic equations look like those in Eqs. (1921) with the Coliolis-like force and the odd-viscosity term. One can derive linearized equations similar to those obtained from active hydrodynamics of polar active matter and numerically confirm its nonzero Chern number. The importance of odd-viscosity terms in the bulk-boundary correspondence in neural progenitor cells has also been discussed [275]. More recently, another group has found chiral flows at the boundary of bacterial colonies [278], which also seems to have a topological origin.